OpenStudy (anonymous):

-5X-Y=25 3X+5Y=7 SOLVE FOR THE FOLLOWING BY ELIMINATING TH E X VARIABLE

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

wait i solved a similar problem of urs,have u checked?

OpenStudy (anonymous):

-5X-Y=25 -5X=25+Y -X=5+Y/5 X=-5-Y/5 subsititing value of X in second equation, 3(-5-Y/5)+5Y=7 -15-3Y/5+5Y=7 -3Y/5+5Y=7+15 (-3Y+25Y)/5=22 22Y/5=22 Y/5=1 Y=5